AAVSOnet status, 2016-03-18

Fri, 03/18/2016 - 11:38

The network is having some problems these days.  I'm hoping most of them will be solved shortly!

BSM_South is operational, but with a loaner CCD camera (an ST-10XME from Peter Nelson).  SBIG is currently repairing the ST-8XME camera, and we expect to return to the original configuration in a couple of weeks.  However, what is a bigger issue right now is that the HQ pipeline processing has been broken for a while due to flatfield file naming.  This issue will be resolved in two phases.  First, all images from January 2016 to present have not been processed at all, and will be processed first.  We're hoping that this will happen in the next week or two.  Second, all images from August 2015 to December 2015 will be reprocessed using better flats.  This may or may not affect your photometry - basically, for that segment of time, old flats were being used and we can do better.

BSM_Hamren has been down for camera repairs.  The system has also been moved to Ocean View Estates, a much drier location than the original Volvano HI.  We hope to have everything working again in a week or two.

BSM_NM has been down due to the CGEM mount failure.  A replacement mount has been sent to Bill Stein, and hopefully will be installed in a week or so.

BSM_HQ is down while we upgrade the software.  This system is our test-bed for upgrades, and it has been a challenge to get ACP Version 8 running.  Once we solve these problems, we'll have a better understanding of what needs to be done for the remaining AAVSOnet systems.

BSM_Berry just had a mount failure.  Greg Bolt is ordering repair parts, but expect it to be down for several weeks.

Coker30 is having problems with its guide camera.  Bill Goff is working on this, but no resolution as of now.  Some data is being taken, and I think we can work around this problem until the camera issue is resolved.

TMO61's dome shutter is still broken.  No word from Jon Holtzman as to when this repair will be done.

So as of today, only BSM_South, SRO50 and OC61 are running nominally!  I'll post when things improve.